Question 1.11[4]


There are two bags of chocolates. Bag one has 5 Barones and 2 KitKats and Bag 2 has 2 Barones


and the 7 KitKats. A chocolate is selected at random from Bag one and added to Bag two. A


chocolate is now drawn randomly from Bag two. Given that the chocolate selected is a KitKat what


is the probability that the original chocolate drawn from Bag one was a Barone? Show all working


out.

1
Expert's answer
2022-06-07T17:38:26-0400

Let "B" denote the event "a Barone is selected at random from Bag one".

Let "K" denote the event "a KitKat is selected at random from Bag one".

Let "k" denote the event "a KitKat is selected at random from Bag two".


"P(B|k)=\dfrac{P(k|B)P(B)}{P(k|B)P(B)+P(k|K)P(K)}"

Given "P(B)=\dfrac{5}{7}, P(K)=\dfrac{2}{7},"


"P(k|B)=\dfrac{7}{2+7+1}=\dfrac{7}{10},"


"P(k|K)=\dfrac{7+1}{2+7+1}=\dfrac{8}{10}"



"P(B|k)=\dfrac{\dfrac{7}{10}(\dfrac{5}{7})}{\dfrac{7}{10}(\dfrac{5}{7})+\dfrac{8}{10}(\dfrac{2}{7})}=\dfrac{35}{51}"